Project

General

Profile

2018-02-20-MEMOMIPO

Joint MICE Experiment Management Office and MICE International Project Office

20th February 2018; 15:30 GMT
RAL, R1 2nd Floor East, MICE Meeting room

Phone details

Present: CR, DR, KL, CW, PG, ABr, AK, JB, MP

  1. Introduction: KL
    • Meeting will capture status of decommissioning and take stock of analysis. Need to review preparation for paper w/s and CM50 which is next week.
  2. Notes of previous meeting and status of actions: All
    • Notes accepted. Status of the actions:
    • CW: Organise test of QD/QP systems for SSs.
      • In hand: System recovered, but, need to reload some of the constants required for the QP/QD. Will excercise the system during the collaboration meeting.
    • CW: Bracing bars. CW to discuss with JB.
      • Underway. Bracing pieces, which have been installed, but, can not rely on them since they place strain on the vacuum vessels. Need to replace the bracing performed by the FC using M12 threaded rod in every second bolt hole. Braces will remain at the end of the solenoids in order to restrain movement.
      • Will purchase helium so that we may be able to recover the magnet.
    • KL/CW: Issue first version of the decommissioning plan this week.
      • Done.
    • CW: Contact Oxford Machine shop re-diffuser.
      • Done. We will remove and send to Oxford post metrology. Also material specifications required too.
  3. Hall status, stability of systems, planning for decommissioning and making safe to be complete by 31Mar18: CW
    • Installation of tracker monuments is now underway. Survey will be done tomorrow. Yellow light will be used during the alignment.
    • Not yet clear whether both trackers will be surveyed this week. Will follow up next week if it is not possible to do both.
    • Once extracted the trackers will be surveyed in RAL metrology.
    • Q7--9 are elecrically disconnected and the water is disconnected too. When mapping of SSU is underway, we shall require to remove Q9 along with its mount.
    • Have not yet removed tracker cryos. Cryo-3 is ready to move, Cryo-4 is almost ready to go. By the end of teh week expect to have both Cryos 1 and 2 ready to move.
    • TOFs 1 & 2, KL and EMR have all been moved.
    • Sumitomo compressors are off, but water is still flowing.
    • FC compressors are off and the magnet is warming to room temperature. Pressure is 1150 mBar. He-level monitor is switched off. Heater control is still running, but, it is not operational.
    • Tomorrow will remove instrumentation, make sure alarm handler is quiet and prepare to remove the module.
    • Most of the compressors are leaking He. Will need to top-up pressure before the magnets are operated.
  4. Data validation ahead of start of Cycle 2017/04: CR
    • Item superceded. Not discussed.
  5. Software and computing status: DR
    • DR has started reviews of the TOF s/w. A first review meeting has taken place. Priority action is to improve the efficiency. This is underway.
    • Tracker simulation mis-handled the calibration which resulted in a poor track-finding efficiency in the MC. This has been fixed. Noise in data is not reproduced in the MC. In data there is evidence that some tracks are being missed. Present thought is that this is a reflection of the low-pt inefficency.
    • DR working on removing issues in the geometrical description of the diffuser. Material descriptio of brass and tungsten is not presently correct. Some runs (LiH, field on) appear to have incorrectracker positions. Geometry update will be available before the reprocessing.
    • New version of MAUS made available to allow data processing on the WWW.
    • VB 7469: CH field systematic simulations have been processed and are ready for analysis.
    • On the issue of decommissioning of the computing in the Hall, DR has discussed with C.~Brew who is supporting the decommissioning.
  6. Memo on measurements required before experiment is decommissioned: CR
    • No update.
    • CR: Will make list of magnet settings for field mapping exercise.
  7. Paper status and plans: CR
    • Emittance measurement paper: will not do a referees meeting next week, but should be one soon after CM50.
    • Scattering paper: J.Nugent will issue an updated MICE note today. Referees meeting on Monday 26Feb18.
    • Emittance evolution: C.Rogers preparing a MICE note for the end of the week. Will not have full systematic analysis. Some issues have been noted, e.g.: momentum vs eloss in MC as mentioned by DR; also cluster-finding inefficiency in the tracker. Possibly other issues to be addressed too. Issue is to be ready for the April VC so that the results can go to IPAC.
    • System performance paper: P.Franchini now helping to guide this forward. CM50 will be a necesary discussion of what will happen.
    • Wedge paper: To be discussed next week.
  8. Discussion of CM50 agenda and issues to be addressed: All
    • CW: hardware decommissioning messages for CM50:
      • Get tracker position measurement completes the post-run re-survey. Remaining issue is the measurement of the top-hat He windows. These will be measured in situ and compared to drawing. To remove these the pieces will have to be sectioned.
      • Mapping of SSD/SSU will then take place starting 19-23Mar18.
      • Between 23Mar18 and 31Mar18 the magnets will be warmed up and the equipment will be put into a quiescent state. Seeking to make a plan for the recovery of equipment that can be used for future experiments.
      • Planning some analysis w/s planned for April. Need to prepare travel approvals ahead of the financial year.
      • Repatriation of equipment: plan needs to be made at the CM50.
      • PSAG on 06Mar18: need to present a clear plan for the removal of equipment in order to provoke the discussion of recovering some resource to be recycled into the analysis.
    • DR: All services and equipment in RR1 and RR2 will be moved to PPD. Some transfer of function has already been executed. P.Franchini, H.Nebrensky and DR are working through the hardware and service move requirements.
    • CR: Need to agree publication plan for next 6 months at CM50.
  9. AOB
    • None.

Summary of actions:
  • CW: Organise test of QD/QP systems for SSs.
  • CW: Bracing bars. CW to discuss with JB.
  • CR: Will make list of magnet settings for field mapping exercise.

Back to MEMO:

Back

Back to MIPO:

Back

Documents